Let A = {x ∈ R : x 2 = 2} and B = {x ∈ R : x ≥ 0}.


1. Find A ∩ B.


2. Find A ∪ B.


3. Find A − B.


4. For U = R, find Ac and Bc .


5. Find N − B.





1
Expert's answer
2021-12-17T13:24:01-0500

1.

"A\cap B=\{\sqrt{2}\}"


2.

"A\cup B=\{x ∈ R : x ≥ 0\ or\ x=-\sqrt{2}\}"

3.

"A- B=\{-\sqrt{2}\}"

4.


"A^C=\{x ∈ R : x^ 2 \not= 2\}"

"B^C=\{x ∈ R : x<0\}"

5.

The set of natural numbers


"\N = \{0, 1, 2, 3, . . .\}"


"\N-B=\empty"